The length of a rectangle is 5 more than twice the width the perimeter is 130 inches what is the area
Answer by JulietG(1812)

2L + 2W = 130
L = 2W+5
2(2W+5) + 2W = 130
4W+10 + 2W = 130
6W + 10 = 130
6W = 120
W = 20
If the width is 20, then the length is twice that plus 5, or 45
L = 45 and W = 20
area = L*W or 900